keke12:

1. (01.02 MC) The number line shows the distance in meters of two divers, A and B, from a shipwreck located at point X: a horizontal number line extends from negative 3 to positive 3. The point labeled as A is at negative 2.5, the point 0 is labeled as X, and the point labeled B is at 1.5 Part A: Write an expression using subtraction to find the distance between the two divers. (5 points) Part B: Show your work and solve for the distance using additive inverses. (5 points)

jhonyy9:

try draw it

toga:

Part A: The expression using subtraction to find the distance between the two divers is B - A. Part B: Distance = B - A Distance = 1.5 - (-2.5) Distance = 1.5 + 2.5 Distance = 4 meters